# Dynamical Low-Rank Compression of Neural Networks with Robustness under Adversarial Attacks

### v1

@@ A.4 Cifar10
A.5 Computational hardware
@@
Appendix B Proofs
@@
$\begin{split}{\mathcal{R}}(S)^{2}&=(S^{\top}\!S-\alpha_{S}^{2}I,S^{\top}\!S-\alpha_{S}^{2}I)\\ &=\|S^{\top}\!S\|^{2}-2\alpha_{S}^{2}(S^{\top}\!S,I)+\alpha_{S}^{4}(I,I)\\ &=\|S^{\top}\!S\|^{2}-\tfrac{1}{r}\|S\|^{4}\\ &=\sum_{i=1}^{r}\sigma_{i}(S^{\top}\!S)^{2}-\frac{1}{r}\bigg{(}\sum_{i=1}^{r}\sigma_{i}(S)^{2}\bigg{)}^{2}\\ &=r\bigg{(}\frac{1}{r}\sum_{i=1}^{r}\sigma_{i}(S^{\top}\!S)^{2}-\bigg{(}\frac{1}{r}\sum_{i=1}^{r}\sigma_{i}(S)^{2}\bigg{)}^{2}\bigg{)}\end{split}$ (26)
Since $S^{\top}\!S$ is symmetric positive semi-definite, $\sigma_{i}(S^{\top}\!S)=\sigma_{i}(S)^{2}$. Applying this substitution yields (3). The proof is complete. ∎
Proof of 1.
@@
Proof of 2.
From (3) and noting $\sigma_{i}(S^{\top}\!S)=\sigma_{i}(S)^{2}$, we obtain
$\frac{1}{r}{\mathcal{R}}(S)^{2}=\frac{1}{r}\sum_{i=1}^{r}\sigma_{i}(S^{\top}\!S)^{2}-\bigg{(}\frac{1}{r}\sum_{i=1}^{r}\sigma_{i}(S^{\top}\!S)\bigg{)}^{2}.$ (31)
From (31), $\tfrac{1}{r}{\mathcal{R}}(S)^{2}$ is the variance of the sequence $\{\sigma_{i}(S^{\top}\!S)\}_{i=1}^{r}$. The Von Szokefalvi Nagy inequality [27] bounds the variance of a finite sequence of numbers below by the range of the sequence (see [34]). Applied to (31), this yields
$\frac{1}{r}{\mathcal{R}}(S)^{2}\geq\frac{(\sigma_{1}(S^{\top}\!S)-\sigma_{r}(S^{\top}\!S))^{2}}{2r}=\frac{(\sigma_{1}(S)^{2}-\sigma_{r}(S)^{2})^{2}}{2r}.$ (32)
@@
$\sqrt{2}{\mathcal{R}}(S)\geq\sigma_{1}(S)^{2}-\sigma_{r}(S)^{2}.$ (33)
An application of the Mean Value Theorem for logarithms (see [28, Proof of Theorem 2.2]), gives
$\ln(\kappa(S))\leq\frac{\sigma_{1}(S)^{2}-\sigma_{r}(S)^{2}}{2\sigma_{r}(S)^{2}}.$ (34)
@@
$\ln(\kappa(S))\leq\frac{1}{\sqrt{2}\sigma_{r}(S)^{2}}{\mathcal{R}}(S),$ (35)
@@
Testing (36) by $S-W$ and rearranging yields
@@
[…]

### v2

@@ B.4 Cifar10
B.5 ImageNet-1k
The ImageNet dataset consists of 1000 classes and over 1.2 million RGB training images, with a standard resolution of $224\times 224$ pixels. We follow the standard data augmentation pipeline for ImageNet, which includes a random resized crop to $224\times 224$, and normalization using mean $[0.5,0.5,0.5]$ and standard deviation $[0.5,0.5,0.5]$. The test set is only resized and center-cropped to $224\times 224$, followed by normalization. Adversarial attacks are generated on the normalized, resized images.
B.6 Computational hardware
@@
Appendix C Proofs
To facilitate the proofs, we remark the definition of L-continuity: A function $f(x)$ is Lipschitz continuous on a domain $D$ if there exists a constant $L\geq 0$ such that for all $x,y\in D$,
$\|f(x)-f(y)\|\leq L\|x-y\|.$
The smallest such $L$ is called the Lipschitz constant.
@@ Appendix C Proofs
$\begin{split}{\mathcal{R}}(S)^{2}&=(S^{\top}\!S-\alpha_{S}^{2}I,S^{\top}\!S-\alpha_{S}^{2}I)\\ &=\|S^{\top}\!S\|^{2}-2\alpha_{S}^{2}(S^{\top}\!S,I)+\alpha_{S}^{4}(I,I)\\ &=\|S^{\top}\!S\|^{2}-\tfrac{1}{r}\|S\|^{4}\\ &=\sum_{i=1}^{r}\varsigma_{i}(S^{\top}\!S)^{2}-\frac{1}{r}\bigg(\sum_{i=1}^{r}\varsigma_{i}(S)^{2}\bigg)^{2}\\ &=r\bigg(\frac{1}{r}\sum_{i=1}^{r}\varsigma_{i}(S^{\top}\!S)^{2}-\bigg(\frac{1}{r}\sum_{i=1}^{r}\varsigma_{i}(S)^{2}\bigg)^{2}\bigg)\end{split}$ (26)
Since $S^{\top}\!S$ is symmetric positive semi-definite, $\varsigma_{i}(S^{\top}\!S)=\varsigma_{i}(S)^{2}$. Applying this substitution yields (3). The proof is complete. ∎
Proof of ˜1.
@@
Proof of ˜2.
From (3) and noting $\varsigma_{i}(S^{\top}\!S)=\varsigma_{i}(S)^{2}$, we obtain
$\frac{1}{r}{\mathcal{R}}(S)^{2}=\frac{1}{r}\sum_{i=1}^{r}\varsigma_{i}(S^{\top}\!S)^{2}-\bigg(\frac{1}{r}\sum_{i=1}^{r}\varsigma_{i}(S^{\top}\!S)\bigg)^{2}.$ (31)
From (31), $\tfrac{1}{r}{\mathcal{R}}(S)^{2}$ is the variance of the sequence $\{\varsigma_{i}(S^{\top}\!S)\}_{i=1}^{r}$. The Von Szokefalvi Nagy inequality [33] bounds the variance of a finite sequence of numbers below by the range of the sequence (see [41]). Applied to (31), this yields
$\frac{1}{r}{\mathcal{R}}(S)^{2}\geq\frac{(\varsigma_{1}(S^{\top}\!S)-\varsigma_{r}(S^{\top}\!S))^{2}}{2r}=\frac{(\varsigma_{1}(S)^{2}-\varsigma_{r}(S)^{2})^{2}}{2r}.$ (32)
@@
$\sqrt{2}{\mathcal{R}}(S)\geq\varsigma_{1}(S)^{2}-\varsigma_{r}(S)^{2}.$ (33)
An application of the Mean Value Theorem for logarithms (see [34, Proof of Theorem 2.2]), gives
$\ln(\kappa(S))\leq\frac{\varsigma_{1}(S)^{2}-\varsigma_{r}(S)^{2}}{2\varsigma_{r}(S)^{2}}.$ (34)
@@
$\ln(\kappa(S))\leq\frac{1}{\sqrt{2}\varsigma_{r}(S)^{2}}{\mathcal{R}}(S),$ (35)
@@
Testing (36) by SWS-W and rearranging yields
@@
Using Hölder’s inequality, the sub-multiplicative property of $\|\cdot\|$, and Young’s inequality, we bound the right hand side of (37) by